OpenStudy (anonymous):

can someone explain how to multiply ((9x-3)/(3x+21)) / ((x+7)/(6x-2))

OpenStudy (anonymous):

I would factor everything you can first. (3x+21) = 3(x+7) (9x-3) = 3(3x-1) (6x-2) = 2(3x-1)

OpenStudy (anonymous):

Then stuff will start canceling

myininaya (myininaya):

\[\frac{\frac{9x-3}{3x+21}}{\frac{x+7}{6x-2}}*\frac{\frac{6x-2}{x+7}}{\frac{6x-2}{x+7}}\] \[=\frac{9x-3}{3x+21} *\frac{6x-2}{x+7}\]

OpenStudy (anonymous):

(9x-3)(6x-2)/(3x+21)(x+7) 6(3x-1)(3x-1)/3(x+7)(x+7) 2(3x-1)^2/(x+7)^2 2((3x-1)/(x+7))^2

OpenStudy (anonymous):

\[2(3x-1)^{2}\div(x-7)^{2}\]

OpenStudy (anonymous):

If you want a simplification, multiply top and bottom by (3x+21), then do the same with (6x-2) \[(9x-3)/((x+7)(3x+21)/(6x-2))\] \[(9x-3)(6x-2)/(3x ^{2}+42x+147)\] \[(54x ^{2}-36x+6)/(3x ^{2}+42x+147)\] You can then factorise top and bottom: \[6(3x-1) ^{2}/3(x+7) ^{2}\] Which gives \[2(3x-1) ^{2}/(x+7) ^{2}\]

OpenStudy (anonymous):

\[\large \frac{\frac{9x-3}{3x+21}}{\frac{x+7}{6x-2}} = \frac{\frac{3(3x-1)}{3(x+7)}}{\frac{x+7}{2(3x-1)}} = \frac{\frac{3x-1}{x+7}}{\frac{x+7}{2(3x-1)}} \cdot \frac{2(x+7)(3x-1)}{2(x+7)(3x-1)} \]\[\large = \frac{2(3x-1)^2}{(x+7)^2} = 2(\frac{3x-1}{x+7})^2\]

myininaya (myininaya):

show off!

myininaya (myininaya):

i thought you were totally against people doing all the work, polpak

OpenStudy (anonymous):

Always when dealing with complex fractions: To convert the complex fraction into a standard fraction multiply the top and bottom by the LCM of all the inner fractions.

OpenStudy (anonymous):

4 other people had already done the work. I was just doing it to show an alternative method.

myininaya (myininaya):

you don't delete stuff anymore?

OpenStudy (anonymous):

And I'm more against people giving answers than I am people giving all the work. Though I think walking them through it is the best way.

OpenStudy (anonymous):

I only delete things if an answer is given with no justification etc. And even then I cannot be everywhere so stuff gets by anyway.

myininaya (myininaya):

yeah i always work it out for them if i don't walk them through it

myininaya (myininaya):

peace i gotta go